OpenStudy (anonymous):

Find the slope of the graph of the function at the given point. Function: g(t)=2+3 cos t Point: (pi, -1)

OpenStudy (anonymous):

@DDCamp

OpenStudy (ddcamp):

First step is to find the derivative at t = pi. What's the derivative of 2+3cos(t)?

OpenStudy (anonymous):

its -3sin(t)

OpenStudy (ddcamp):

OK. The tangent line is the line that passes through a graph, but has the same slope of the graph at that point. What would the slope of g(x) be at t=pi?

OpenStudy (anonymous):

Hmmm...I have no idea

OpenStudy (anonymous):

-3sin/pi?

OpenStudy (anonymous):

@DDCamp

OpenStudy (ddcamp):

Yes, -3sin(pi) sin(pi) = 0, so the slope of the tangent line is 0 at (pi, -1)

OpenStudy (ddcamp):

Now that you know the slope, and you have a point, you can plug this into point-slope form: y-y₁ = m(x-x₁) y-(-1) = 0(x-pi) y+1 = 0 y = -1

OpenStudy (ddcamp):

Sorry, I have to go now. Good luck with the rest of your work.

OpenStudy (anonymous):

Wow thank you! So my answer would be y=-1?

OpenStudy (anonymous):

Or is that just the check?